OpenStudy (anonymous):

7 and 5 sixths-2 and 5 ninths=???

OpenStudy (anonymous):

\[\left(7+\frac{5}{6}\right)-\left(2+\frac{5}{9}\right)=\frac{95}{18}=5\frac{5}{18} \]